# Structured and Balanced Multi-component and Multi-layer Networks

### v1

@@ 3.2 High dimensional cases
Let us now consider the extension to multi-dimension using two dimensions as an example since the simple dimension-by-dimension strategy applies to any dimension.
@@
$\begin{cases}{\mathcal{L}}_{1,i}(a_{i})=x_{i-1},\\ {\mathcal{L}}_{1,i}(b_{i})=x_{i}\end{cases}\quad\textnormal{and}\quad\,\ \begin{cases}{\mathcal{L}}_{2,j}(c_{i})=y_{j-1},\\ {\mathcal{L}}_{2,j}(d_{i})=y_{j}.\end{cases}$ (5)
@@
$\left\{\begin{array}[]{l}f_{i,0}(x,y)\coloneqq f\Big{(}{\mathcal{L}}_{1,i}(x),\,y\Big{)},\\ f_{0,j}(x,y)\coloneqq f\Big{(}x,\,{\mathcal{L}}_{2,j}(y)\Big{)},\\ f_{i,j}(x,y)\coloneqq f\Big{(}{\mathcal{L}}_{1,i}(x),\,{\mathcal{L}}_{2,j}(y)\Big{)}=f_{0,j}\Big{(}{\mathcal{L}}_{1,i}(x),\,y\Big{)}=f_{i,0}\Big{(}x,\,{\mathcal{L}}_{2,j}(y)\Big{)}.\end{array}\right.$ (6)
@@
$\psi_{i}(x)=\begin{cases}b_{i}&\textnormal{if }x>x_{i},\\ {\mathcal{L}}_{1,i}^{-1}(x)&\textnormal{if }x\in[x_{i-1},x_{i}],\\ a_{i}&\textnormal{if }x<x_{i-1}\end{cases}\quad\textnormal{and}\quad\phi_{j}(y)=\begin{cases}d_{j}&\textnormal{if }y>y_{j},\\ {\mathcal{L}}_{2,j}^{-1}(y)&\textnormal{if }y\in[y_{j-1},y_{j}],\\ c_{j}&\textnormal{if }y<y_{j-1}.\end{cases}$ (7)
The following result provides a decomposition of $f$ that fits into the structure of MMNN.
@@
$\begin{split}f(x,\,y)&=\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}\Big{(}\psi_{i}(x),\,\phi_{j}(y)\Big{)}-\sum_{i=1}^{n}\sum_{j=1}^{m-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j}\Big{)}\\ &\quad\ -\sum_{i=1}^{n-1}\sum_{j=1}^{m}f_{0,j}\Big{(}x_{i},\,\phi_{j}(y)\Big{)}+\sum_{i=1}^{n-1}\sum_{j=1}^{m-1}f(x_{i},\,y_{j})\end{split}$ (8)
@@
$\begin{split}f(x,\,y)=\sum_{i=1}^{n}f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}-\sum_{i=1}^{n-1}f(x_{i},\,y).\end{split}$
For each $i$, using the 1D decomposition technique described in Section 3.1, we find the decompositions for $f_{i,0}\big{(}\psi_{i}(x),\,y\big{)}$ and $f(x_{i},\,y)$. We have
$\begin{split}&\phantom{=\;}\sum_{j=1}^{m}f_{i,j}\Big{(}\psi_{i}(x),\,\phi_{j}(y)\Big{)}=\sum_{j=1}^{m}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}\circ\phi_{j}(y)\Big{)}\\ &=\sum_{j=1}^{\ell-1}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}\circ\phi_{j}(y)\Big{)}+f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,\ell}\circ\phi_{\ell}(y)\Big{)}+\sum_{j=\ell+1}^{m}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}(\circ\phi_{j}(y))\Big{)}\\ &=\sum_{j=1}^{\ell-1}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}(d_{j})\Big{)}+f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,\ell}\circ{\mathcal{L}}_{2,\ell}^{-1}(y)\Big{)}+\sum_{j=\ell+1}^{m}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}(c_{j})\Big{)}\\ &=\sum_{j=1}^{\ell-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j}\Big{)}+f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}+\sum_{j=\ell+1}^{m}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j-1}\Big{)}\\ &=f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}+\sum_{j=1}^{m-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j-1}\Big{)},\end{split}$
@@
$\begin{split}f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}=\sum_{j=1}^{m}f_{i,j}\Big{(}\psi_{i}(x),\,\phi_{j}(y)\Big{)}-\sum_{j=1}^{m-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j}\Big{)}.\end{split}$
@@
$\begin{split}f(x_{i},\,y)=\sum_{j=1}^{m}f_{0,j}\Big{(}x_{i},\,\phi_{j}(y)\Big{)}-\sum_{j=1}^{m-1}f(x_{i},\,y_{j}).\end{split}$
Therefore, for any $(x,y)\in[x_{k-1},x_{k}]\times[y_{\ell-1},y_{\ell}]$,
$\begin{split}&\phantom{=\;\;}f(x,\,y)=\sum_{i=1}^{n}f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}-\sum_{i=1}^{n-1}f(x_{i},\,y)\\ &=\sum_{i=1}^{n}\Bigg{(}\sum_{j=1}^{m}f_{i,j}\Big{(}\psi_{i}(x),\,\phi_{j}(y)\Big{)}-\sum_{j=1}^{m-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j}\Big{)}\Bigg{)}-\sum_{i=1}^{n-1}\Bigg{(}\sum_{j=1}^{m}f_{0,j}\Big{(}x_{i},\,\phi_{j}(y)\Big{)}-\sum_{j=1}^{m-1}f(x_{i},\,y_{j})\Bigg{)}\\ &=\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}\Big{(}\psi_{i}(x),\,\phi_{j}(y)\Big{)}-\sum_{i=1}^{n}\sum_{j=1}^{m-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j}\Big{)}-\sum_{i=1}^{n-1}\sum_{j=1}^{m}f_{0,j}\Big{(}x_{i},\,\phi_{j}(y)\Big{)}+\sum_{i=1}^{n-1}\sum_{j=1}^{m-1}f(x_{i},\,y_{j}).\end{split}$
Since $k$ and $\ell$ are arbitrary, the above equation holds for all $(x,y)=\cup_{k=1}^{n}\cup_{\ell=1}^{m}[x_{k-1},x_{k}]\times[y_{\ell-1},y_{\ell}]=[x_{0},x_{n}]\times[y_{0},y_{m}]$. ∎
@@ 3.3 Related work
The KA representation theorem [15] states that any multivariate continuous function on a hypercube can be expressed as a finite composition of continuous univariate functions and the binary operation of addition. However, this elegant mathematical representation may result in non-smooth or even fractal univariate functions in general due to this very specific form of representation, a computational challenge one has to address in practice. KA representation has been explored in several studies [17, 20, 32, 13]. A recently proposed network known as the KA network (KAN) utilizes spline functions to approximate the univariate functions in the KA representation. The proposed MMNN is motivated by a multi-component and multi-layer smooth decomposition, or a “divide and conquer” approach, employing distinct network architectures, activation functions, and training strategies.
@@ 4 Numerical experiments
In this section, we perform extensive experiments to validate our analysis and demonstrate the effectiveness of MMNNs through multi-component and multi-layer decomposition studied in Section 3. In particular, our tests show its ability in 1) adaptively capturing localized high-frequency features in Section 4.1, 2) approximating highly oscillatory functions in Section 4.2, and 3) extending to higher dimensions in Section 4.3 as well as some interesting learning dynamics in Section 4.4. All our experiments involve target functions that include high-frequency components in various ways and are difficult to handle by shallow networks (no matter how wide) as shown in our previous work [38]. Moreover, our experience on these tests shows that using a fully connected deep neural network would require many more parameters and is much harder (if possible) to train to get a comparable result. This is mainly due to a balanced and structured network design of MMNN in terms of 1) the network width $w$, which is the number of hidden neurons or random basis functions in each component, 2) the rank $r$, which is the number of components in each layer, and 3) the network depth $l$, which is the number of layers in the network. The use of a controllable number of collective components (through ${\bm{A}}$) in each layer instead of a large number of individual neurons and the use of fixed and randomly chosen weights $({\bm{W}},{\bm{b}})$ make the training process more effective.
In all tests, 1) data are sampled enough to resolve fine features in the target function, 2) the Adam optimizer is used in training, 3) mean squared error (MSE) is the loss function, 4) all parameters are initialized according to the PyTorch default initialization (see Section 2.3) unless otherwise specified, 5) ${\bm{W}}$’s and ${\bm{b}}$’s (the parameters inside the activation functions, see Section 2.2) are fixed and only ${\bm{A}}$’s and ${\bm{c}}$’s (the parameters outside the activation functions) are trained, 6) computations are conducted on an NVIDIA RTX 3500 Ada Generation Laptop GPU (power cap 130W), with most experiments concluding within a range from a few dozen to several thousand seconds. All our MMNN setups are specified by three parameters $(w,r,l)$ which depends on the function complexity. Another tuning parameter is the learning rate the choice of which is guided by the following criteria: 1) not too large initially due to stability, 2) a decreasing rate with iterations such that the learning rate becomes small near the equilibrium to achieve a good accuracy while not decreasing too fast (especially during a long training process for more difficult target functions) so that the training is stalled.
@@ 4.1 Localized rapid changes
We begin with two examples in 1D. The first is $f(x)=\arctan(100x+20)$, which is smooth but features a rapid transition at zero. While demonstrated in our previous work [38], a shallow network struggles to capture such a simple local fast transition which contains high-frequencies, we show that this function can be approximated easily by a composition of a smooth function on top of a (repeated) spatial decomposition and local rescaling using MMNN structure in Section 2.2. Our test indeed verifies that our new architecture can effectively capture a localized fast transition rather easily using a very small network of size $(16,4,3)$ as shown in Figure 11. For this test, a total of 1000 data points are uniformly sampled in the range $[-1,1]$, with a mini-batch size of 100, a learning rate of $10^{-3}$, and the number of epochs set to 2000. Figure 11 gives the error plot.
@@
(a) FEM.
(b) MMNN.
Figure 14: (a) Least square using equally spaced 153 FEM bases. (b) MMNN with $(16+1)\times 4\times(3-1)+(16+1)=153$ free parameters.
Now we show an example in 2D shown in Figure 16 and defined in polar coordinates by
$f(r,\theta)=\begin{cases}0&\textnormal{if }0.5+25\rho-25r\leq 0,\\ 1&\textnormal{if }0.5+25\rho-25r\geq 1,\\ 0.5+25\rho-25r&\textnormal{otherwise},\\ \end{cases}\quad\textnormal{where}\quad\rho=0.1+0.02\cos(8\pi\theta).$
Again a rather compact MMNN of size $(100,10,6)$ can produce a good approximation. Figure 17 shows the error during the training process and Figure 16 shows the $\log$ plot of training and testing errors in MSE. For this test there are a total of $400^{2}$ uniformly sampled points in $[-1,1]^{2}$ with mini-batch size of $1000$ and a learning rate of $10^{-3}\times 0.9^{\lfloor k/25\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},1000$ is the epoch number. We compare the result with piecewise linear interpolation and least square approximation using FEM basis on a uniform grid with the same number of degrees of freedom in Figure 18. As observed before, MMNN renders the best result due to its adaptivity through training.
@@
Figure 16: Training and test errors measured in MSE.
(a) Epoch 10.
(b) Epoch 50.
(c) Epoch 100.
(d) Epoch 1000.
Figure 17: The error during the training process.
@@
(d) Network (difference).
(e) Interpolation (difference).
(f) FEM (difference).

### v2

@@ 4.2 High dimensional cases
Let us now consider the extension to multiple dimensions, using the case of two dimensions as an example, since the straightforward dimension-by-dimension strategy can be applied to any number of dimensions.
@@
(7) $\begin{cases}{\mathcal{L}}_{1,i}(a_{i})=x_{i-1},\\ {\mathcal{L}}_{1,i}(b_{i})=x_{i}\end{cases}\quad\textnormal{and}\quad\,\ \begin{cases}{\mathcal{L}}_{2,j}(c_{i})=y_{j-1},\\ {\mathcal{L}}_{2,j}(d_{i})=y_{j}.\end{cases}$
@@
(8) $\left\{\begin{array}[]{l}f_{i,0}(x,y)\coloneqq f\Big{(}{\mathcal{L}}_{1,i}(x),\,y\Big{)},\\ f_{0,j}(x,y)\coloneqq f\Big{(}x,\,{\mathcal{L}}_{2,j}(y)\Big{)},\\ f_{i,j}(x,y)\coloneqq f\Big{(}{\mathcal{L}}_{1,i}(x),\,{\mathcal{L}}_{2,j}(y)\Big{)}=f_{0,j}\Big{(}{\mathcal{L}}_{1,i}(x),\,y\Big{)}=f_{i,0}\Big{(}x,\,{\mathcal{L}}_{2,j}(y)\Big{)}.\end{array}\right.$
@@
(9) $\psi_{i}(x)=\begin{cases}b_{i}&\textnormal{if }x>x_{i},\\ {\mathcal{L}}_{1,i}^{-1}(x)&\textnormal{if }x\in[x_{i-1},x_{i}],\\ a_{i}&\textnormal{if }x<x_{i-1}\end{cases}\quad\textnormal{and}\quad\phi_{j}(y)=\begin{cases}d_{j}&\textnormal{if }y>y_{j},\\ {\mathcal{L}}_{2,j}^{-1}(y)&\textnormal{if }y\in[y_{j-1},y_{j}],\\ c_{j}&\textnormal{if }y<y_{j-1}.\end{cases}$
The theorem below provides a decomposition of $f$ that fits into the MMNN structure.
@@
(10) $\begin{split}f(x,\,y)&=\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}\Big{(}\psi_{i}(x),\,\phi_{j}(y)\Big{)}-\sum_{i=1}^{n}\sum_{j=1}^{m-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j}\Big{)}\\ &\quad\ -\sum_{i=1}^{n-1}\sum_{j=1}^{m}f_{0,j}\Big{(}x_{i},\,\phi_{j}(y)\Big{)}+\sum_{i=1}^{n-1}\sum_{j=1}^{m-1}f(x_{i},\,y_{j})\end{split}$
@@
$f(x,\,y)=\sum_{i=1}^{n}f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}-\sum_{i=1}^{n-1}f(x_{i},\,y).$
For each $i$, using the one-dimensional decomposition technique described in Section 4.1, we find the decompositions for $f_{i,0}\big{(}\psi_{i}(x),\,y\big{)}$ and $f(x_{i},\,y)$. We have
$\begin{split}&\phantom{=\;}\sum_{j=1}^{m}f_{i,j}\Big{(}\psi_{i}(x),\,\phi_{j}(y)\Big{)}=\sum_{j=1}^{m}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}\circ\phi_{j}(y)\Big{)}\\ &=\sum_{j=1}^{\ell-1}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}\circ\phi_{j}(y)\Big{)}+f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,\ell}\circ\phi_{\ell}(y)\Big{)}+\sum_{j=\ell+1}^{m}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}\circ\phi_{j}(y)\Big{)}\\ &=\sum_{j=1}^{\ell-1}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}(d_{j})\Big{)}+f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,\ell}\circ{\mathcal{L}}_{2,\ell}^{-1}(y)\Big{)}+\sum_{j=\ell+1}^{m}f_{i,0}\Big{(}\psi_{i}(x),\,{\mathcal{L}}_{2,j}(c_{j})\Big{)}\\ &=\sum_{j=1}^{\ell-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j}\Big{)}+f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}+\sum_{j=\ell+1}^{m}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j-1}\Big{)}\\ &=f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}+\sum_{j=1}^{m-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j-1}\Big{)},\end{split}$
@@
(11) $\begin{split}f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}=\sum_{j=1}^{m}f_{i,j}\Big{(}\psi_{i}(x),\,\phi_{j}(y)\Big{)}-\sum_{j=1}^{m-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j}\Big{)}.\end{split}$
@@
(12) $\begin{split}f(x_{i},\,y)=\sum_{j=1}^{m}f_{0,j}\Big{(}x_{i},\,\phi_{j}(y)\Big{)}-\sum_{j=1}^{m-1}f(x_{i},\,y_{j}).\end{split}$
Therefore, for any $(x,y)\in[x_{k-1},x_{k}]\times[y_{\ell-1},y_{\ell}]$, by (11) and (12), we have
$\begin{split}f(x,\,y)&=\sum_{i=1}^{n}f_{i,0}\Big{(}\psi_{i}(x),\,y\Big{)}-\sum_{i=1}^{n-1}f(x_{i},\,y)\\ &=\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}\Big{(}\psi_{i}(x),\,\phi_{j}(y)\Big{)}-\sum_{i=1}^{n}\sum_{j=1}^{m-1}f_{i,0}\Big{(}\psi_{i}(x),\,y_{j}\Big{)}\\ &\quad\ -\sum_{i=1}^{n-1}\sum_{j=1}^{m}f_{0,j}\Big{(}x_{i},\,\phi_{j}(y)\Big{)}+\sum_{i=1}^{n-1}\sum_{j=1}^{m-1}f(x_{i},\,y_{j}).\end{split}$
Since $k$ and $\ell$ are arbitrary, the above equation holds for all $(x,y)=\cup_{k=1}^{n}\cup_{\ell=1}^{m}[x_{k-1},x_{k}]\times[y_{\ell-1},y_{\ell}]=[x_{0},x_{n}]\times[y_{0},y_{m}]$.
@@ 4.3 Related work
Several lines of research are closely related to this work, including approximation theory, low-rank methods, random feature models, and architectures inspired by the Kolmogorov–Arnold representation.
@@
The KA representation theorem [17] states that any multivariate continuous function on a hypercube can be expressed as a finite composition of continuous univariate functions and the binary operation of addition. However, this elegant mathematical representation may result in compositions of non-smooth or even fractal univariate functions in general, a computational challenge one has to address in practice. KA representation has been explored in several studies [19, 22, 35, 14]. A recently proposed network known as the KA network (KAN) utilizes spline functions to approximate the univariate functions in the KA representation. The proposed MMNN is motivated by a multi-component and multi-layer smooth decomposition, or a “divide and conquer” approach, employing distinct network architectures, activation functions, and training strategies.
@@ 5 Numerical experiments
We perform extensive experiments to validate our analysis and demonstrate the effectiveness of MMNNs through multi-component and multi-layer decomposition studied in Section 4. In particular, our tests show its ability in 1) adaptively capturing localized high-frequency features in Section 5.1, 2) approximating highly oscillatory functions in Section 5.2, and 3) some interesting learning dynamics in Section 5.4. All our experiments involve target functions that include high-frequency components in various ways and are difficult to handle by shallow networks (no matter how wide) as shown in our previous work [41]. Moreover, our experience on these tests shows that using a fully connected deep neural network would require many more parameters and is much harder (if possible) to train to get a comparable result. This is mainly due to a balanced and structured network design of MMNN in terms of 1) the network width $w$, which is the number of hidden neurons or random basis functions in each component, 2) the rank $r$, which is the number of components in each layer, and 3) the network depth $l$, which is the number of layers in the network. The use of a controllable number of collective components (through ${\bm{A}}$) in each layer instead of a large number of individual neurons and the use of fixed and randomly chosen weights $({\bm{W}},{\bm{b}})$ make the training process more effective.
In all tests, 1) data are sampled enough to resolve fine features in the target function, 2) the Adam optimizer is used in training, 3) the mean squared error (MSE) is the loss function, 4) all computation and training use single precision in PyTorch, 5) all parameters are initialized according to the PyTorch default initialization (Section 2.2) unless otherwise specified, 6) ${\bm{W}}$’s and ${\bm{b}}$’s (the parameters inside the activation functions, see Section 2.1) are fixed and only ${\bm{A}}$’s and ${\bm{c}}$’s (the parameters outside the activation functions) are trained, 7) computations are conducted on a NVIDIA RTX 3500 Ada Generation Laptop GPU (power cap 130W), with most experiments concluding within a range from a few dozen to several thousand seconds. All our MMNN setups are specified by three parameters $(w,r,l)$ which depend on the function complexity. Another tuning parameter is the learning rate which is guided by the following criteria: 1) not too large initially due to stability, 2) a decreasing rate with iterations such that the learning rate becomes small near the equilibrium to achieve a good accuracy while not decreasing too fast (especially during a long training process for more difficult target functions) so that the training is stalled.
@@ 5.1 Localized rapid changes
We begin with two one-dimensional examples. The first is $f(x)=\arctan(100x+20)$, which is smooth but features a rapid transition at zero. While demonstrated in our previous work [41], a shallow network struggles to capture such a simple local fast transition which contains high-frequencies, we show that this function can be approximated easily by a composition of a smooth function on top of a (repeated) spatial decomposition and local rescaling using MMNN structure in Section 2.1. Our test indeed verifies that our new architecture can effectively capture a localized fast transition rather easily using a very small network of size $(16,4,3)$ as shown in Figure 13. For this test, a total of 1000 data points are uniformly sampled in the range $[-1,1]$, with a mini-batch size of 100, a learning rate of $10^{-3}$, and the number of epochs set to 2000. Figure 13 gives the error plot.
@@
Figure 16: Left: Least square using equally spaced 153 FEM bases. Right: MMNN with $(16+1)\times 4\times(3-1)+(16+1)=153$ free parameters.
Finally, we show an example in the two-dimensional case shown in Figure 18 and defined in polar coordinates by
$f(r,\theta)=\begin{cases}0&\textnormal{if }0.5+25\rho-25r\leq 0,\\ 1&\textnormal{if }0.5+25\rho-25r\geq 1,\\ 0.5+25\rho-25r&\textnormal{otherwise},\\ \end{cases}\quad\textnormal{where}\ \rho=0.1+0.02\cos(8\pi\theta).$
Again a rather compact MMNN of size $(100,10,6)$ can produce a good approximation. Figure 18 shows the $\log$ plot of training and testing errors in MSE. For this test there are a total of $400^{2}$ uniformly sampled points in $[-1,1]^{2}$ with mini-batch size of $1000$ and a learning rate of $10^{-3}\times 0.9^{\lfloor k/25\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},1000$ is the epoch number. We compare the result with piecewise linear interpolation and least square approximation using FEM basis on a uniform grid with the same number of degrees of freedom in Figure 19. As observed before, MMNN renders the best result due to its adaptivity through training. When adaptive finite element is applicable, it is hard to beat. However, adaptive FEMs are more humanly involved, while MMNNs based on training are more automatic. The key message here is to demonstrate that MMNN has adaptive features through training/optimization, which will be useful when an adaptive finite element method becomes difficult or impossible in applications.
@@
Figure 18: Errors (MSE).
@@
(d) Network (error).
(e) Interpolation (error).
(f) FEM (error).